Fifteen (15) years ago Jennifer was quarter of the age she would be fifteen (15) years form now. How old is Jennifer?: Fifteen (15) years ago Jennifer was quarter of the age she would be fifteen (15) years form now. How old is Jennifer?
Answer by stanbon(18792)

Fifteen (15) years ago Jennifer was quarter of the age she would be fifteen (15) years form now. How old is Jennifer?
--------------
Let Jennifer's age 15 years ago be x.
Fifteen years from now she will be x+30
--------------
EQUATION:
x = (1/4)(x+30)
4x = x+30
3x = 30
x = 10 (Jennifer's age 15 years ago)
10+15 = 25 (Jernnifer's age now)
=================
Cheers,
Stan H.